# Instance-Optimal Imprecise Convex Hull

### v3

@@ 2.1 Explicit dynamic planar convex hull
We rely on a data structure to maintain the upper quarter convex hull of a dynamic point set. Overmars and van Leeuwen [20] provide the best-known solution, achieving deterministic update time $O(\log^{2}n)$. Their data structure, known as the Partial Hull Tree (PHT), is now textbook material. We follow their terminology:
Definition 2.6 (PHT from [20]).
@@ 3 Witnesses and an instance-optimal reconstruction strategy
We introduce a new geometric classification over the edges of $\operatorname{OCH}(F)$. This leads to an alternative instance-optimal reconstruction strategy, detailed in Algorithm 1. Our results hold in full generality for any family $F$ of points and closed regions whose boundaries are simple, closed, piecewise-$C^{1}$ curves. In particular, each region is connected and bounded, and any non-point region has a connected interior equal to its closure; see Figure 2. We define a classification over vertex pairs $(s,t)$, which we apply to edges of $\operatorname{OCH}(F)$. However, we formulate the classification for arbitrary vertex pairs, as this generality will be useful for our data structures. We say that two regions $R_{a},R_{b}\in F$ are strictly vertically separated if there exists a vertical line $\ell$ where $R_{a}$ and $R_{b}$ lie in opposite open half-planes defined by $\ell$.
@@
Algorithm 1 A data structure friendly reconstruction strategy: Reconstruct$(F)$.
@@
In the following two subsections, we first show that each case in Algorithm 1 retrieves witness. Then, we prove that if Algorithm 1 is instance-optimal.
3.1 Proving that Algorithm 1 retrieves only witnesses
The first case of Algorithm 1 considers a non-canonical edge, which implies that there is a point on $\operatorname{OCH}(F)$ that lies in two regions that are not both point regions.
@@ 3.1 Proving that Algorithm 1 retrieves only witnesses
[…]
The second case of Algorithm 1 considers an edge $(s,t)$ of $\operatorname{OCH}(F)$ that is canonical but not dividing in $F$. Let $s\in V(R_{a})$ and $t\in V(R_{b})$. Because $(s,t)$ is canonical but not dividing it holds that $a\neq b$ and $R_{a}$ and $R_{b}$ are distinct regions that are not strictly vertically separated. The following lemma implies that $\{R_{a},R_{b}\}$ is indeed a witness.
@@
Let $P^{\prime}\sim(F-R_{a}-R_{b})$ be arbitrary. Since $R_{a}$ and $R_{b}$ appear on $\operatorname{OCH}(F)$, there are vertices $q,r$ on $\operatorname{OCH}(F)$ with $q\in R_{a}$ and $r\in R_{b}$. If $q=r$, Lemma 3.4 shows that $\{R_{a},R_{b}\}$ is a witness. Otherwise, set $p_{a}=q,p_{b}=r$, then $R_{a}$ and $R_{b}$ appear at distinct points on $\operatorname{CH}(P^{\prime}\cup\{p_{a},p_{b}\})$. On the other hand, as $R_{a}$ and $R_{b}$ are not vertically separated, there exist $s\in R_{a},t\in R_{b}$ with the same $x$-coordinate. Set $p_{a}^{\prime}=s,p_{b}^{\prime}=t$, then $R_{a}$ and $R_{b}$ cannot appear at distinct points on $\operatorname{CH}(P^{\prime}\cup\{p_{a}^{\prime},p_{b}^{\prime}\})$. Indeed, if $s=t$, then $R_{a}$ and $R_{b}$ form duplicate points, and if $s\neq t$, then at most one of $R_{a}$ and $R_{b}$ appears on $\operatorname{CH}(P^{\prime}\cup\{p_{a}^{\prime},p_{b}^{\prime}\})$. This shows that $P^{\prime}\cup\{R_{a},R_{b}\}$ is not finished.
The third case of Algorithm 1 considers an occupied edge. The next lemma shows that any choice of region that has a vertex in the occupied band gives rise to a witness.
@@
The final case of Algorithm 1 considers any remaining spanning subchains of $\operatorname{OCH}(F)$.
@@
3.2 Proving that Algorithm 1 is instance-optimal
We now prove that Algorithm 1 is instance-optimal. That is, the algorithm retrieves the minimal number of regions (up to constant factors) necessary to determine the ordering of points on the convex hull, for any realisation $P\sim F$. We first observe that Algorithm 1 only terminates when $F$ is terminal, meaning: every edge of $\operatorname{OCH}(F)$ is dividing, no edge of $\operatorname{OCH}(F)$ is occupied, and no contiguous subchain of $\operatorname{OCH}(F)$ is spanning.
@@ 3.2 Proving that Algorithm 1 is instance-optimal
With Lemma 3.12, we define for each region $R$ that appears on $\operatorname{OCH}(F)$ its left (or right) neighbour as the first region on $\operatorname{OCH}(F)$ that lies strictly left (or right) of $R$. Note that two of duplicate point regions on $\operatorname{OCH}(F)$ have the same neighbours. Next, we show that the points corresponding to regions in $A$ always appear on the convex hull (Lemma 3.14), and that no points corresponding to regions in $F-A$ appear on the convex hull (Lemma 3.16).
@@
Let $P^{\prime}\sim A$ and $R_{b}\in A-\{(-\infty,-\infty),(\infty,-\infty)\}$ be arbitrary. Let $R_{a}$ and $R_{c}$ be the left and right neighbour of $R_{b}$ on $\operatorname{OCH}(F)$. Observe that these neighbours are well-defined as (,)($\infty$,$\infty$)( - ∞ , - ∞ ) and (,)($\infty$,$\infty$)( ∞ , - ∞ ) lie on $\operatorname{OCH}(F)$.
@@
Let $P^{\prime}\sim A$ be arbitrary. Suppose by contradiction that there is a region $R_{i}\in F-A$ with a point $p_{i}\in R_{i}$ on or outside of $\operatorname{CH}(P^{\prime})$. Since the inside of $\operatorname{CH}(P^{\prime})$ is convex, we may require $p_{i}$ to be an extreme point of $R_{i}$, so $p_{i}\in V(R_{i})$. By Lemma 3.12 and Lemma 3.14, there exist consecutive regions $R_{a},R_{b}\in A$ for which $p_{i}$ lies on or above $p_{a}p_{b}$. As $p_{a}p_{b}$ lies fully inside $\operatorname{band}(R_{a},R_{b})$, the point $p_{i}$ lies in or directly above $\operatorname{band}(R_{a},R_{b})$. We make a case distinction, where $p_{i}$ to lies directly above $R_{b}$ (or $R_{a}$), or, on the upper tangent of $R_{a},R_{b}$.
@@
Algorithm 1 is an instance-optimal reconstruction strategy.
@@
First, we show that when Algorithm 1 terminates, the set $F$ is finished. Let $F$ be terminal. Let $A\subseteq F$ be the multiset of regions that appear on $\operatorname{OCH}(F)$. Let $P^{\prime}\sim A$ and $Q^{\prime}\sim F-A$ be arbitrary. By Lemma 3.16, no points in $Q^{\prime}$ are on $\operatorname{CH}(P^{\prime}\cup Q^{\prime})$. By Lemma 3.14, the points in $P^{\prime}$ all appear on $\operatorname{CH}(P^{\prime}\cup Q^{\prime})$. Moreover, together with Lemma 3.12, this implies that for all $P_{1}\sim A$ and $P_{2}\sim A$, $\operatorname{\preceq}(\operatorname{CH}(P_{1}))=\operatorname{\preceq}(\operatorname{CH}(P_{2}))$. So, $F$ is finished. Hence, Algorithm 1 is a correct reconstruction strategy.
[…]
@@ 4 Regions that are simple $k$-gons
Figure 8: The convex chain from $q$ to $v$ is hit in $F$ due to the red vertices.
@@
[by Definition 3.2] For any canonical bridge $(s,t)$ in $T$, there exists a unique area $R(s)$ that contains $s$. $R(s)$ is either a point, or a unique region in $F$ and can be found in $O(1)$ time by checking one of the leafs that $(s,t)$ points to.
@@
Next, consider candidate chains of length $\geq 4$ in which $(s,t)$ is an interior edge. Check, in $O(1)$ time via Observation 4.1, whether both $s$ and $t$ lie in a unique region $R_{b}$. If so, perform an in-order traversal of $\mathbb{E}(\nu)$—moving up to $O(k)$ steps left from $s$ and up to $O(k)$ steps right from $t$—to find the maximal contiguous chain with all edges dividing and interior vertices contained in $V(R_{b})$. The chain ends when we encounter an edge whose endpoints do not lie entirely in $V(R_{b})$. This identifies the unique candidate chain $C$ of length $\geq 4$ containing $(s,t)$ as an interior edge. To find chains where $(s,t)$ is the first or last edge, apply the above procedure to the edges immediately before and after $(s,t)$.
@@
Test, using Observation 4, whether any bridge is canonical in constant time. Since the recourse is $O(k\log m)$, the set of all bridges can be updated without overhead.
To maintain $\Lambda^{*}(\nu)$ for each node $\nu\in T$, we follow the technique from [20] for concatenable queues. For any node $\nu$ with child $x$, compute $\Lambda(x)$ from $(\Lambda(\nu),e(\nu),\Lambda^{*}(x))$ in $O(\log m)$ time by splitting $\Lambda(\nu)$ at the bridge $e(\nu)$, and joining the result with $\Lambda^{*}(x)$. Traverse all $O(k)$ updated root-to-leaf paths, using the split operation, in $O(k\log^{2}m)$ total time. Then, traverse the paths bottom-up. At each node $\nu$ with children $x$ and $y$, we have access to the newly updated trees $\Lambda(x)$ and $\Lambda(y)$. Compute $(\Lambda(\nu),\Lambda^{*}(x),\Lambda^{*}(y))$ by splitting $\Lambda(x)$ and $\Lambda(y)$ at the new bridge $e(\nu)$ and joining the result.
@@
Apply Lemma 4.3 to obtain all $O(1)$ candidate chains in $\mathbb{E}(\nu)$ containing $(s,t)$. Let $C=(q,\ldots,r^{\prime})$ be such a chain. To test whether $C$ is spanning in $F$, use Observation 4.1 to find the unique regions $R(q)$, $R_{b}$, and $R(r^{\prime})$. Construct $\operatorname{OCH}(R(q)\cup R(r^{\prime}))$ in $O(k\log k)$ time and perform a convex hull intersection test with $R_{b}$ in $O(\log m)$ time using the algorithm of Chazelle [7]. The chain is spanning if and only if this test returns true.
@@
Apply Lemma 4.3 to find all candidate chains containing $(x,y)$. Then, for each candidate chain, consider the adjacent edge and apply the lemma again to construct all possible hit chains. For each resulting subchain, test whether the middle edge i$(s,t)$ is occupied as follows: Construct $\operatorname{band}(R(s),R(t))$ in $O(k\log k)$ time [8]. Temporarily remove $R(s)$ and $R(t)$ from $F$, and delete $s$ and $t$ from $V(F)$, in $O(k\log^{2}m)$ time. Update the corresponding PHT $T^{\prime}$ without updating auxiliary data. The hull $\operatorname{CH}(V(F)\setminus\{s,t\})$ is stored at the root of $T^{\prime}$. By intersection testing between $\operatorname{band}(R(s),R(t))$ and $\operatorname{CH}(T^{\prime})$ in $O(\log m)$ time [7], we can test if any hit chain is occupied.
@@
As with Corollary 4.11, we invoke Lemma 4.13 on $O(k\log m)$ nodes where the leaf set changes. For each, we update the associated trees using split and join operations.
@@
In the next section, we note that our approach has two bottlenecks. First, the algorithm from Lemma 4.3 may ‘skip’ over $O(k)$ edges to find the largest candidate chain. By applying a technique similar to skip-lists, we improve its running time by a factor $k$. Secondly, our approach frequently uses a subroutine where for any two regions $R(s)$ and $R(t)$ we construct $\operatorname{band}(R(s),R(t))$ in $O(k\log k)$ time. We show that by storing for each region $R_{i}$ its convex hull, this construction time becomes polylogarithmic and we obtain the following:

### v4

@@ 2.1 Explicit dynamic planar convex hull
We rely on a data structure to maintain the upper quarter convex hull of a dynamic point set. Overmars and van Leeuwen [DBLP:conf/stoc/OvermarsV80] provide the best-known solution, achieving deterministic update time $O(\log^{2}n)$. Their data structure, known as the Partial Hull Tree (PHT), is now textbook material. We follow their terminology:
Definition 2.6 (PHT from [DBLP:conf/stoc/OvermarsV80]).
@@ 3 Witnesses and an instance-optimal reconstruction strategy
We introduce a new geometric classification over the edges of $\operatorname{OCH}(F)$. This leads to an alternative instance-optimal reconstruction strategy, detailed in Algorithm 1. Our results hold in full generality for any family $F$ of points and closed regions whose boundaries are simple, closed, piecewise-$C^{1}$ curves. In particular, each region is connected and bounded, and any non-point region has a connected interior equal to its closure; see Figure˜2. We define a classification over vertex pairs $(s,t)$, which we apply to edges of $\operatorname{OCH}(F)$. However, we formulate the classification for arbitrary vertex pairs, as this generality will be useful for our data structures. We say that two regions $R_{a},R_{b}\in F$ are strictly vertically separated if there exists a vertical line $\ell$ where $R_{a}$ and $R_{b}$ lie in opposite open half-planes defined by $\ell$.
@@
Algorithm 1 A data structure friendly reconstruction strategy: Reconstruct(F)(F).
@@
In the following two subsections, we first show that each case in Algorithm˜1 retrieves a witness. Then, we prove that if Algorithm˜1 is instance-optimal.
3.1 Proving that Algorithm˜1 retrieves only witnesses
The first case of Algorithm˜1 considers a non-canonical edge, which implies that there is a point on $\operatorname{OCH}(F)$ that lies in two regions that are not both point regions.
@@ 3.1 Proving that Algorithm˜1 retrieves only witnesses
Let $P^{\prime}\sim(F-R_{a}-R_{b})$ be arbitrary. Since $R_{a}$ and $R_{b}$ appear on $\operatorname{OCH}(F)$, there are vertices $q,r$ on $\operatorname{OCH}(F)$ with $q\in R_{a}$ and $r\in R_{b}$. If $q=r$, Lemma˜3.4 shows that $\{R_{a},R_{b}\}$ is a witness. Otherwise, set $p_{a}=q,p_{b}=r$, then $R_{a}$ and $R_{b}$ appear at distinct points on $\operatorname{CH}(P^{\prime}\cup\{p_{a},p_{b}\})$. On the other hand, as $R_{a}$ and $R_{b}$ are not vertically separated, there exist $s\in R_{a},t\in R_{b}$ with the same $x$-coordinate. Set $p_{a}^{\prime}=s,p_{b}^{\prime}=t$, then $R_{a}$ and $R_{b}$ cannot appear at distinct points on $\operatorname{CH}(P^{\prime}\cup\{p_{a}^{\prime},p_{b}^{\prime}\})$. Indeed, if $s=t$, then $R_{a}$ and $R_{b}$ form duplicate points, and if $s\neq t$, then at most one of $R_{a}$ and $R_{b}$ appears on $\operatorname{CH}(P^{\prime}\cup\{p_{a}^{\prime},p_{b}^{\prime}\})$. This shows that $P^{\prime}\cup\{R_{a},R_{b}\}$ is not finished.
The second case of Algorithm˜1 considers an edge $(s,t)$ of $\operatorname{OCH}(F)$ that is canonical but not dividing in $F$. Let $s\in V(R_{a})$ and $t\in V(R_{b})$. Because $(s,t)$ is canonical but not dividing it holds that $a\neq b$ and $R_{a}$ and $R_{b}$ are distinct regions that are not strictly vertically separated. The following lemma implies that $\{R_{a},R_{b}\}$ is indeed a witness.
@@
Let $P^{\prime}\sim(F-R_{a}-R_{b})$ be arbitrary. Since $R_{a}$ and $R_{b}$ appear on $\operatorname{OCH}(F)$, there are vertices $q,r$ on $\operatorname{OCH}(F)$ with $q\in R_{a}$ and $r\in R_{b}$. If $q=r$, Lemma˜3.4 shows that $\{R_{a},R_{b}\}$ is a witness. Otherwise, set $p_{a}=q,p_{b}=r$, then $R_{a}$ and $R_{b}$ appear at distinct points on $\operatorname{CH}(P^{\prime}\cup\{p_{a},p_{b}\})$. On the other hand, as $R_{a}$ and $R_{b}$ are not vertically separated, there exist $s\in R_{a},t\in R_{b}$ with the same $x$-coordinate. Set $p_{a}^{\prime}=s,p_{b}^{\prime}=t$, then $R_{a}$ and $R_{b}$ cannot appear at distinct points on $\operatorname{CH}(P^{\prime}\cup\{p_{a}^{\prime},p_{b}^{\prime}\})$. Indeed, if $s=t$, then $R_{a}$ and $R_{b}$ form duplicate points, and if $s\neq t$, then at most one of $R_{a}$ and $R_{b}$ appears on $\operatorname{CH}(P^{\prime}\cup\{p_{a}^{\prime},p_{b}^{\prime}\})$. This shows that $P^{\prime}\cup\{R_{a},R_{b}\}$ is not finished.
The third case of Algorithm˜1 considers an occupied edge. The next lemma shows that any choice of region that has a vertex in the occupied band gives rise to a witness.
@@
The final case of Algorithm˜1 considers any remaining spanning subchains of $\operatorname{OCH}(F)$.
@@
3.2 Proving that Algorithm˜1 is instance-optimal
We now prove that Algorithm˜1 is instance-optimal. That is, the algorithm retrieves the minimal number of regions (up to constant factors) necessary to determine the ordering of points on the convex hull, for any realisation $P\sim F$. We first observe that Algorithm˜1 only terminates when $F$ is terminal, meaning: every edge of $\operatorname{OCH}(F)$ is dividing, no edge of $\operatorname{OCH}(F)$ is occupied, and no contiguous subchain of $\operatorname{OCH}(F)$ is spanning.
@@ 3.2 Proving that Algorithm˜1 is instance-optimal
With Lemma 3.12, we define for each region $R$ that appears on $\operatorname{OCH}(F)$ its left (or right) neighbour as the first region on $\operatorname{OCH}(F)$ that lies strictly left (or right) of $R$. Note that two duplicate point regions on $\operatorname{OCH}(F)$ have the same neighbours. Next, we show that the points corresponding to regions in $A$ always appear on the convex hull (Lemma 3.14), and that no points corresponding to regions in FAF-A appear on the convex hull (Lemma 3.16).
@@
Let $P^{\prime}\sim A$ and $R_{b}\in A-\{(-\infty,-\infty),(\infty,-\infty)\}$ be arbitrary. Let $R_{a}$ and $R_{c}$ be the left and right neighbour of $R_{b}$ on $\operatorname{OCH}(F)$. Observe that these neighbours are well-defined as $(-\infty,-\infty)$ and $(\infty,-\infty)$ lie on $\operatorname{OCH}(F)$.
@@
Let $P^{\prime}\sim A$ be arbitrary. Suppose by contradiction that there is a region $R_{i}\in F-A$ with a point $p_{i}\in R_{i}$ on or outside of $\operatorname{CH}(P^{\prime})$. Since the inside of $\operatorname{CH}(P^{\prime})$ is convex, we may require $p_{i}$ to be an extreme point of $R_{i}$, so $p_{i}\in V(R_{i})$. By Lemma 3.12 and Lemma 3.14, there exist consecutive regions $R_{a},R_{b}\in A$ for which $p_{i}$ lies on or above $p_{a}p_{b}$. As $p_{a}p_{b}$ lies fully inside $\operatorname{band}(R_{a},R_{b})$, the point $p_{i}$ lies in or directly above $\operatorname{band}(R_{a},R_{b})$. We make a case distinction, where $p_{i}$ lies directly above $R_{b}$ (or $R_{a}$), or, on the upper tangent of $R_{a},R_{b}$.
@@
Algorithm˜1 is an instance-optimal reconstruction strategy.
@@
First, we show that when Algorithm˜1 terminates, the set $F$ is finished. Let $F$ be terminal. Let $A\subseteq F$ be the multiset of regions that appear on $\operatorname{OCH}(F)$. Let $P^{\prime}\sim A$ and $Q^{\prime}\sim F-A$ be arbitrary. By Lemma˜3.16, no points in $Q^{\prime}$ are on $\operatorname{CH}(P^{\prime}\cup Q^{\prime})$. By Lemma 3.14, the points in $P^{\prime}$ all appear on $\operatorname{CH}(P^{\prime}\cup Q^{\prime})$. Moreover, together with Lemma 3.12, this implies that for all $P_{1}\sim A$ and $P_{2}\sim A$, $\operatorname{\preceq}(\operatorname{CH}(P_{1}))=\operatorname{\preceq}(\operatorname{CH}(P_{2}))$. So, $F$ is finished. Hence, Algorithm˜1 is a correct reconstruction strategy.
We now show instance-optimality. By Lemmas 3.4, 3.6, 3.8 and 3.10, the algorithm retrieves a constant-size witness in each iteration. Let $X_{1},\dots,X_{k}$ be the sets of non-point regions of these witnesses. Since the witnesses have constant size, Algorithm˜1 does $O(k)$ retrievals. A retrieval turns a region into a point region, hence the $X_{i}$ are pairwise disjoint. On the other hand, if $(F\hskip-1.42271pt\odot_{\hskip-1.42271pt\textit{P}}\hskip-0.85355ptA)$ is finished for some $A\subseteq F$, then $A$ needs to contain at least one element from each $X_{i}$ (by the definition of witness). As the $X_{i}$ are disjoint, this forces $|A|\geq k$. Hence, Algorithm˜1 is instance optimal by Observation 2.
@@ 4 Regions that are simple $k$-gons
Figure 8: The convex chain from $q$ to $v$ is hit in $F$ since $(s,t)$ is occupied due to the red vertices.
@@
[by Definition 3.2] For any canonical bridge $(s,t)$ in $T$, there exists a unique area $R(s)$ that contains $s$. $R(s)$ is either a point, or a unique region in $F$ and can be found in $O(1)$ time by checking the regions lists stored at one of the leaves that $(s,t)$ points to.
@@
Next, consider candidate chains of length $\geq 4$ in which $(s,t)$ is an interior edge. Check, in $O(1)$ time via Observation 4.1, whether both $s$ and $t$ lie in a unique region $R_{b}$. If so, perform an in-order traversal of $\mathbb{E}(\nu)$—moving up to $O(k)$ steps left from $s$ and up to $O(k)$ steps right from tt—to find the maximal contiguous chain with all edges dividing and interior vertices contained in $V(R_{b})$. The chain ends when we encounter an edge whose endpoints do not lie entirely in $V(R_{b})$. This identifies the unique candidate chain $C$ of length $\geq 4$ containing $(s,t)$ as an interior edge. To find chains where $(s,t)$ is the first or last edge, apply the above procedure to the edges immediately before and after $(s,t)$.
@@
Test, using Observation 4, whether any bridge is canonical in constant time. Since the recourse is $O(k\log m)$, the set of all bridges can be updated without incurring asymptotic overhead.
To maintain $\Lambda^{*}(\nu)$ for each node $\nu\in T$, we follow the technique from [DBLP:conf/stoc/OvermarsV80] for concatenable queues. For any node $\nu$ with child $x$, compute $\Lambda(x)$ from $(\Lambda(\nu),e(\nu),\Lambda^{*}(x))$ in $O(\log m)$ time by splitting $\Lambda(\nu)$ at the bridge $e(\nu)$, and joining the result with $\Lambda^{*}(x)$. Traverse all $O(k)$ updated root-to-leaf paths, using the split operation, in $O(k\log^{2}m)$ total time. Then, traverse the paths bottom-up. At each node $\nu$ with children $x$ and $y$, we have access to the newly updated trees $\Lambda(x)$ and $\Lambda(y)$. Compute $(\Lambda(\nu),\Lambda^{*}(x),\Lambda^{*}(y))$ by splitting $\Lambda(x)$ and $\Lambda(y)$ at the endpoints of the new bridge $e(\nu)$ and joining the result.
@@
Apply Lemma 4.3 to obtain all $O(1)$ candidate chains in $\mathbb{E}(\nu)$ containing $(s,t)$. Let $C=(q,\ldots,r^{\prime})$ be such a chain. To test whether $C$ is spanning in $F$, use Observation 4.1 to find the unique regions $R(q)$, $R_{b}$, and $R(r^{\prime})$. Construct $\operatorname{OCH}(R(q)\cup R(r^{\prime}))$ in $O(k\log k)$ time and perform a convex hull intersection test with $R_{b}$ in $O(\log m)$ time using the algorithm of Chazelle [chazelle1980detection]. The chain is spanning if and only if this test returns true.
@@
Apply Lemma 4.3 to find all candidate chains containing $(x,y)$. Then, for each candidate chain, consider the adjacent edge and apply the lemma again to construct all possible hit chains. For each resulting subchain, test whether the middle edge i(s,t)(s,t) is occupied as follows: Construct $\operatorname{band}(R(s),R(t))$ in $O(k\log k)$ time [de2008computational]. Temporarily remove $R(s)$ and $R(t)$ from $F$, and delete $s$ and $t$ from $V(F)$, in $O(k\log^{2}m)$ time. Update the corresponding PHT $T^{\prime}$ without updating auxiliary data. The hull $\operatorname{CH}(V(F)\setminus\{s,t\})$ is stored at the root of $T^{\prime}$. By intersection testing between $\operatorname{band}(R(s),R(t))$ and $\operatorname{CH}(T^{\prime})$ in $O(\log m)$ time [chazelle1980detection], we can test if any hit chain is occupied.
@@
As with Corollary 4.11, we invoke Lemma˜4.13 on $O(k\log m)$ nodes where the leaf set changes. For each, we update the associated trees using split and join operations.
@@
In the full version, we note that our approach has two bottlenecks. First, the algorithm from Lemma 4.3 may ‘skip’ over $O(k)$ edges to find the largest candidate chain. By applying a technique similar to skip-lists, we improve its running time by a factor $k$. Secondly, our approach frequently uses a subroutine where for any two regions $R(s)$ and $R(t)$ we construct $\operatorname{band}(R(s),R(t))$ in $O(k\log k)$ time. We show that by storing for each region $R_{i}$ its convex hull, this construction time becomes polylogarithmic and we obtain the following: